OpenStudy (anonymous):

L lies between K and M on a straight line.KM=14,KL=x+2,and LM=2x . find KL

OpenStudy (anonymous):

KL+LM+KM (x+2)+2x=14 3x=12 x=4 then since KL=x+2 KL=4+2=6

OpenStudy (anonymous):

thnx